Suppose the Fed decided to purchase $100 billion worth of government securities in the open market (assume all payments are are directly deposited into or withdrawn from the banking system). What impact would this action have on the economy? Specifically, answer the following questions:

Instructions: Enter your responses as a whole number.

a. How will M1 be affected initially?

No initial change to M1

Increase by $100 billion CORRECT

Not enough information to answer

Decrease by $100 billion

b. By how much will the banking system’s lending capacity increase if the reserve requirement is 20 percent?

$ billion

c. Must interest rates rise or fall to induce investors to utilize this expanded lending capacity?

Fall CORRECT

Rise

d. By how much will aggregate demand initially increase if investors borrow and spend all the newly available credit?

$ 400 billion CORRECT

e. Under what circumstances would the Fed be pursuing such an open market policy?

Inflation

Recession CORRECT

f. To attain those same objectives, what should the Fed do with the

(i) Discount rate?

Decrease CORRECT

Increase

(ii) Reserve requirement?

Decrease CORRECT

Increase
Answer #1

answer :

a)

Increase by $100 billion

Explanation: M1 will rise by $10 billion, on an assumption that the sellers of the securities hold the proceeds as cash or deposit it in a transactions account.

b)

$80 billion

Explanation: 80 billion Money multiplier = 1 / required reserve ratio = 1/0.20 = 5 Lending capacity = 0.80 * $100 billion = 80 billion  

c)

Fall

Explanation: Interest rates should fall to entice the investors to avail of the expanded lending capacity. Lower interest rates indicate smaller amounts of interest payments on loans with banks.

d)

$400 billion

Explanation: Money multiplier = 1 / required reserve ratio = 1/0.20 = 5 Increase in aggregate demand = lending capacity * money multiplier = 5* 80 billion = 400 billion

e)

FED would pursue recession with a target to stimulate the economy by increasing aggregate demand

f) 1)

Decrease

Explanation: To increase aggregate demand, the Fed would lower the discount rate

2)

Decrease

Explanation: To increase aggregate demand, the Fed would decrease the discount rate